A capacitor charged from a 50V DC supply is discharged across a charge-measuring instrument and found to have carried a charge of 10uC. What is the capacitance of the capacitor and how much energy was stored in it?
1
Expert's answer
2021-02-06T10:06:03-0500

a) We can find the capacitance of the capacitor from the formula:


"Q=CV,""C=\dfrac{Q}{V}=\dfrac{10\cdot10^{-6}\ C}{50\ V}=0.2\cdot10^{-6}\ F=0.2\ \mu F."

b) We can find the energy stored in the capacitor from the formula:


"E=\dfrac{1}{2}CV^2=\dfrac{1}{2}\cdot0.2\cdot10^{-6}\ F\cdot(50\ V)^2=2.5\cdot10^{-4}\ J."